I am having a hard time setting a word problem up. The word problem is this - Carl bought two plots of land for a total of $120,000. When he sold the first plot, he made a profit of 15%. When he sold the second, he lost 10%. His total profit was $5500. How much did he pay for each piece of land?
Found 2 solutions by checkley77, Greenfinch:
Answer by checkley77(12844)

.15x-.10(120,000-x)=5,500
.15x-12,000+.10x=5,500
.25x=5,500+12,000
.25x=17,500
x=17,500/.25
x=70,000 original price of the profit plot of land.
120,000-70,000=50,000
Proof:
.15*70,000-10*50,000=5,500
10,500-5000=5,500
5,500=5,500

Answer by Greenfinch(383)

1st piece of land cost a and made profit of 0.15 a
2nd cost (120000 - a ) and made loss of (120000 - a) 0.1
Adding up 0.15a - 12000 + 0.1a = to a profit of 5500
Net cost is 0.25a - 12000 = 5500
So 0.25 a = 17,500
So a = 70000
Check 15% of 70 000 is 10,500 profit
10% 0f the rest (50000) is 5000 loss
Difference is $5500